OpenStudy (anonymous):

Find the rectangular coordinates of the point with the polar coordinates (3, 3/2 pi).

OpenStudy (anonymous):

(0, 3) (0, -3) (-3, 0) (3, 0)

OpenStudy (ipwnbunnies):

Let (X,Y) be the rectangular coordinate. And let \[(r, \theta)\] be the polar coordinate. X = r*cos(theta) Y = r*sin(theta)

OpenStudy (anonymous):

can you walk me through it? I don't really understand

OpenStudy (ipwnbunnies):

Ok, let's look at the given coordinate. (3, 3pi/2) This means a point exists 3 units away from the origin (0,0) in the direction of the angle 3pi/2.